OpenStudy (anonymous):

could someone help me?how to solve it. gived function R(x) = X² + 2hx + h², so R(x+h) - R(x) / h = ....

OpenStudy (anonymous):

do you mean [ R(x+h) - R(x)] / h OR R(x+h) -[ R(x) / h] ??? ;)

OpenStudy (anonymous):

Sorry I don't mean like that. it's like I write before.

OpenStudy (anonymous):

:D okaY if its [ R(x+h) - R(x)] / h then R(x) = X² + 2hx + h²=(x+h)²[R(x+h) - R(x) ]/ h =[(x+h+h)² - (x+h)²]/h =(x+2h+x+h)(x+2h-x-h)/h =2x+3h

OpenStudy (anonymous):

hope this helps:)